OpenStudy (anonymous):

find: integral (4x^1/5 + 5/x^5)dx

OpenStudy (anonymous):

\[\LARGE \int 4x^{1/5}+\frac{5}{x^5}dx\] \[\LARGE \int 4x^{1/5}dx +\int \frac{5}{x^5}dx\] \[\LARGE 4\int x^{1/5}dx +5\int \frac{1}{x^5}dx\] \[\LARGE 4\int x^{1/5}dx +5\int x^{-5 }dx\] ... can you finish it?

OpenStudy (lgbasallote):

have you tried distributing the integral and solving from there? i.e. \(\LARGE \int 4x^{\frac{1}{5}}dx + \int 5x^{-5}dx\)

OpenStudy (anonymous):

i am trying to finish it give me a sec

OpenStudy (anonymous):

\[\LARGE \int x^rdx=\frac{x^{r+1}}{r+1}+C\]

OpenStudy (anonymous):

is the answer 10/3x^6/5 + 5/6x^6 +C?

OpenStudy (lgbasallote):

-5 + 1 is NOT -6 :D

OpenStudy (anonymous):

would it be 10x^2/5 - 5/4x^4 +C? @lgbasallote

OpenStudy (lgbasallote):

the first term seems abt right

OpenStudy (lgbasallote):

i meant the first term the first time

OpenStudy (anonymous):

OH okay so its 10/3x^6/5 - 5/4x^4 +C

OpenStudy (lgbasallote):

if you mean \(\LARGE \frac{10 x^{\frac{6}{5}}}{3} - \frac{5}{4x^4} + C\) then it's right...your spacings are just confusing hahaha

OpenStudy (anonymous):

thank you so much for the great help @lgbasallote